OpenStudy (anonymous):

Limits Q: The limit of a constant as x goes to infinity. Is the answer the constant or zero?

OpenStudy (anonymous):

constant

OpenStudy (anonymous):

its like saying you have B what is B as A goes to infinity Well A is not even in there so it's just B

OpenStudy (zarkon):

\[\lim_{x\to\infty}c=c\]

OpenStudy (anonymous):

\[\lim (x \to \infty) (2-4x-8x ^{2}) = -\infty\]

OpenStudy (anonymous):

shouldnt the answer then be 2 in that question then?

OpenStudy (zarkon):

no

OpenStudy (zarkon):

\[2-4x-8x^2\] is not a constant function

OpenStudy (anonymous):

ooh i think i get it haha

OpenStudy (anonymous):

thanks!

OpenStudy (zarkon):

yw